# MPCM-Net: Multi-scale network integrates partial attention convolution with Mamba for ground-based cloud image segmentation

### v1

@@ IV-E Ablation Study
To validate the efficacy of key components in our proposed MPCM-Net, comprehensive ablation studies were conducted. Built upon an encoder-decoder architecture, MPCM-Net incorporates the MPAC in the encoder and the Multi Mamba decoding. The MPAC consists of MPC and MPA blocks, primarily comprising multi-scale partial attention modules (ParM, ParSM, and ParAM). The M2B within decoder is fundamentally constructed using SSHD. Therefore, we conduct multiple ablation configurations on the CSRC dataset to verify module effectiveness. Note that MPCM-Net served as the baseline method in all experiments.
[…]
TABLE VII: Comparison experiments of CA and SLA with SE and SA on the CSRC dataset. Baseline is our method, Row1 uses SE instead of CA, Row2 uses SA instead of SLA. $\downarrow$ (or $\uparrow$) indicates lower (or higher) is better. The best results are highlighted in bold.
@@
Figure 11: The segmentation results of the ablation experiments. The results are shown in (a) original images, (b) ground truth, (c) MPCM-Net, (d) MPCM-Net without MPAC, and (e) MPCM-Net without MMD, respectively. The red circles in column (d) indicate the misclassification of scale variable cloud clusters in the results of removing MPAC. The red circles in column (e) indicate the misclassification of the boundary information in the results of removing MMD. Column (c) shows that our method can effectively alleviate these problems.
To prove the proposed modules of the MPCM-Net that can solve the problem of the multi-scale variations from cloud motion across successive time frames and the incomplete boundaries due to the loss of local information, we select some representative segmentation results, as shown in Fig. 11. The cloud clusters in these images include diverse scales and complex edges. The first column (a) and the second column (b) of Fig. 11 are the original images and the corresponding GT, respectively. The third column to the fifth column of Fig. 11 (c) – (e) show the experimental results of the baseline, baseline without the MPAC, and baseline without the MMD, respectively. It should be noted that ResNet-50 is employed as the encoder in the configuration without MPAC. Specifically, from Fig. 11, we can see that the results of our proposed MPCM-Net achieve the closest alignment to the ground truth by comparing the segmentation results of the third column and the fourth to fifth columns, demonstrating that MPCM-Net can effectively improve the segmentation effect of ground-based cloud imagery. The fourth column of Fig. 11 shows that the results of the method without MPAC exhibit inaccurate segmentation across cloud scales. The comparison between columns (c) and (d) shows the critical role of MPAC in multi-scale context extraction. In addition, as shown in column (e), the results of the method without MMD present the degradation of boundary information, attributable to the capacity of MMD for establishing long-range local feature correlations within the decoder.
@@
TABLE VIII: Ablation experiments of MPAC and MMD on the CSRC dataset. $\uparrow$ indicates higher is better. The best results are highlighted in bold.
@@ V Conclusion
Accurate ground-based cloud image segmentation is a critical enabling technology for grid dispatch operations and clean energy integration. In this paper, we addressed fundamental limitations in existing deep learning approaches by proposing MPCM-Net, a novel framework integrating partial convolution in the encoder and Mamba architecture in the decoder. Our principal contributions are threefold: First, we designed a novel MPAC encoder that facilitates interaction between local and global features across diverse cloud scales through ParCM and ParSM. Second, we introduced the SLA mechanism, a high-performance attention variant that delivers multi-scale feature representation with linear complexity. Third, we developed a M2B in the decoder that effectively mitigates contextual information loss while enhancing inference efficiency through spatial-semantic multi-dimensional feature integration. Extensive experimentation on our newly proposed CSRC dataset demonstrates the significant superiority of MPCM-Net over state-of-the-art methods. Comprehensive ablation studies rigorously validate the effectiveness of each module. The results confirm that our MPCM-Net achieves an optimal balance between segmentation accuracy and inference speed for ground-based cloud segmentation. Future work will focus on extending MPCM-Net to ultra-short-term PV power forecasting tasks, further enhancing its practical utility in renewable energy management systems.
[Uncaptioned image] Penghui Niu received the B.S. degree in automobile engineering from Shanghai Dianji University, Shanghai, China, in 2017 and the M.S. degree in instruments science and technology from Tianjin University of Science and Technology, Tianjin, China, in 2021. He is currently pursuing the Ph.D. degree in control science and engineering with the School of Artificial Intelligence, Hebei University of Technology, Tianjin, China. His research interests include ultra-short-term PV power forecasting and intelligent interpretation of remote sensing images.
[Uncaptioned image] Jiashuai She received the B.S. degree in computer science and technology from Hebei University of Technology, Tianjin, China, in 2022. He is currently pursuing the Ph.D. degree in electrical engineering and automation with the School of Electrical Engineering, Hebei University of Technology, Tianjin, China. His research interests include ultra-short-term PV power forecasting and intelligent interpretation of remote sensing images.
[Uncaptioned image] Taotao Cai received the PhD degree from Deakin University, in 2020 after spending two years with the University ofWestern Australia (2017-2019) and over one year at Deakin University (2019-2020) during his Ph.D. studies. He is a lecturer in computing with the University of Southern Queensland (UniSQ). His primary focus is on research and teaching in the field of Data Science, including graph data processing, social network analytics, recommendation systems, and complexity science. Prior to joining the faculty at UniSQ, Taotao held positions as a PostDoctoral Research Fellow at Macquarie University (2021-2023) and an associate research fellow at Deakin University (2020-2021). During this time, he made significant research contributions, which have been published in leading international conferences and journals such as IEEE ICDE, Information Systems, and IEEE TKDE.
[Uncaptioned image] Yajuan Zhang received the B.S. degree in computer science and technology from Inner Mongolia Agricultural University, Inner Mongolia, China, in 2007 and the M.S. degree in control science and engineering from Hebei University of Technology, Tianjin, China, in 2010. She is currently a Senior Experimentalist in the School of Artificial Intelligence at Hebei University of Technology. Her current research interests include image processing and data mining.
[Uncaptioned image] Ping Zhang received the M.S. degree in software engineering from the College of Software, Jilin University, Changchun, China, in 2018, and the Ph.D. degree in computer science from Jilin University, Changchun, in 2021. She is currently a lecturer at the School of Artificial Intelligence, Hebei University of Technology, Tianjin, China. Her research interests include feature selection and machine learning.
[Uncaptioned image] Junhua Gu received the B.S. degree in mathematics from Shanghai Jiaotong University, Shanghai, China, in 1988, and the M.S. degree in computer science and the Ph.D. degree in electrical engineering from the Hebei University of Technology, Tianjin, China, in 1993 and 1997, respectively. He is a Professor with the School of Artificial Intelligence, Hebei University of Technology. He is the Hebei new century “333 Talent Project” second level suitable person. He has published more than 70 papers and the current research interests include big data, intelligent control, and intelligent transportation systems.
[…]

### v2

@@ IV-E Ablation Study
To validate the efficacy of key components in our proposed MPCM-Net, comprehensive ablation studies were conducted. Built upon an encoder-decoder architecture, MPCM-Net incorporates the MPAC in the encoder and the Multi Mamba decoding. The MPAC consists of MPC and MPA blocks, primarily comprising multi-scale partial attention modules (ParM, ParSM, and ParAM). The M2B within the decoder is fundamentally constructed using SSHD. Therefore, we conduct multiple ablation configurations on the CSRC dataset to verify module effectiveness. Note that MPCM-Net served as the baseline method in all experiments.
TABLE VII: Comparison experiments of CA and SLA with SE and SA on the CSRC dataset. Baseline is our method, Row1 uses SE instead of CA, Row2 uses SA instead of SLA.
@@
Figure 10: Visualization results show different categories of the CSRC dataset validation set using Grad-CAM. (a) Input image, (b) MPCM-Net, and (c) MPCM-Net with SE.
[…]
TABLE VIII: Ablation experiments of MPAC and MMD on the CSRC dataset.
@@
Figure 11: The segmentation results of the ablation experiments. The results are shown in (a) original images, (b) ground truth, (c) MPCM-Net, (d) MPCM-Net without MPAC, and (e) MPCM-Net without MMD, respectively. The red rectangles in column (d) indicate the misclassification of scale variable cloud clusters in the results of removing MPAC. The red rectangles in column (e) indicate the misclassification of the boundary information in the results of removing MMD. Column (c) shows that our method can effectively alleviate these problems.
[…]
@@ V Conclusion
The precise ground-based cloud image segmentation constitutes a fundamental prerequisite for grid dispatch operations and clean energy integration. To surmount the inherent limitations in existing DL approaches by proposing MPCM-Net, this article introduces a novel framework integrating partial convolution in the encoder and Mamba architecture in the decoder. Our principal contributions are threefold: First, we designed a novel MPAC encoder that facilitates interaction between local and global features across diverse cloud scales through ParCM and ParSM. Second, we introduced the SLA mechanism, a high-performance attention variant that yields comprehensive multi-scale feature representation. Third, we developed a M2B in the decoder that effectively mitigates contextual information loss through spatial-semantic multi-dimensional feature integration. Extensive experimentation on our newly proposed CSRC dataset demonstrates the significant superiority of MPCM-Net over SOTA methods. Comprehensive ablation studies rigorously validate the effectiveness of each module. The results confirm that our MPCM-Net achieves a superior equilibrium between segmentation accuracy and inference speed for ground-based cloud segmentation. Future work will focus on extending MPCM-Net to ultra-short-term PV power forecasting tasks, further enhancing its practical utility in renewable energy management systems.